# CurvFed: Curvature-Aligned Federated Learning for Fairness without Demographics

### v6

@@ 5.3.1. Feasibility of Real-World Deployment with CurvFed:
Intel
i9-9900K
Processor
Fedavg 1.67 1074.05 0.03 48.12
@@
AMD Ryzen 9
7950X 16-Core
Processor
Fedavg 0.77 1033.62 0.01 24.93
@@
NVIDIA GeForce
RTX 4090
Fedavg 0.14 954.52 0.01 15.0
@@
Google Pixel 6
Fedavg 4.24 370.84 0.12 49.11
@@
Table 5. Training Time (s), Training Process (RSS) Memory Usage (MBs), Inference Time (s), Training Resource usage (%) for FL Approaches across Different Platforms
5.3.2. Communication Overhead:
5.3.3. Resource Utilization and Overhead Breakdown
@@
To contextualize these results, we also benchmarked the upload and download latency for a trained model ($\sim$34.4 MB) across CPU-based platforms, using Google Cloud APIs with PyDrive authentication (Table 7). Communication latency per round is on the order of seconds, which is one to two orders of magnitude higher than the Fisher computation overhead per batch. Since such communication is repeated across many FL rounds (e.g., 80 rounds for WIDAR), network transfer typically dominates the end-to-end runtime (le2024exploring, ).

### v7

@@ Appendix B Runtime benchmarking vs baseline approaches
Intel i9-9900K Processor Fedavg 1.67 1074.05 0.03 48.12
@@
AMD Ryzen 9 7950X 16-Core Processor Fedavg 0.77 1033.62 0.01 24.93
@@
NVIDIA GeForce RTX 4090 Fedavg 0.14 954.52 0.01 15.0
@@
Google Pixel 6 Fedavg 4.24 370.84 0.12 49.11
@@
Table 15. Training Time (s), Training Process (RSS) Memory Usage (MBs), Inference Time (s), Training Resource usage (%) averaged for five runs for FL Approaches across Different Platforms
B.0.1. Resource Utilization and Overhead Breakdown
@@
To contextualize these results, we also benchmarked the upload and download latency for a trained model ($\sim$34.4 MB) across CPU-based platforms, using Google Cloud APIs with PyDrive authentication (Table 17). Communication latency per round is on the order of seconds, which is one to two orders of magnitude higher than the Fisher computation overhead per batch. Since such communication is repeated across many FL rounds (e.g., 80 rounds for WIDAR), network transfer typically dominates the end-to-end runtime (Le et al., 2024).